Two students are selected at random from a class of 13 males and 7 females. Find the probability that both students are males. (See Example 4. Round your answer to two decimal places.)
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(57424)

Two students are selected at random from a class of 13 males and 7 females. Find the probability that both students are males.
----------------------------
# of ways to pick 2 males: 13C2 = 78
-----
# of ways to pick 2 students randomly: 20C2 = 190
-----------------------------------
Probability of selecting 2 males: 78/190 = 0.4105
======================================================
Cheers,
Stan H.

Answer by checkley77(12569)

13/20*12/19=156/420 OR .3714 OR 37.14 % CHANCE BOTH ARE MALES.